if it take one tap 6 mins to fill a bath and another tap 12 mins to fill a bath, how long will it take if both taps are running?
Answer by god2012(113)

Given : time taken by first tab to fill a bath is 6 mins
Therefore, part of the bath filled in 1 min is 1/6 ------------1
Similarly, time taken by second tap is 12 mins
Therefore, part of the bath filled in 1 min is 1/12 ------------2
Part of the bath filled in 1 min when both the taps are open =
part filled by first tap in one min + part filled by 2nd tap in 1 min
= 1/6 + 1/12 = 1/4
Therefore, 1/12 part is filled in 1 min
OR 1 part, that is bath is completely filled in 12 mins when both the taps are on